OpenStudy (anonymous):

Polar coordinates: Find the area of the region enclosed by one loop of the curve. r=1+2sintheta(inner loop) need help finding the limits of integration. I get sintheta=-1/2, now where do i go from there?

OpenStudy (anonymous):

@Hoa still not entirely sure. :l

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i figured it out actually. since sin theta=-1/2, i have to use the quadrants 3 and 4. i was using 1 and 2.

OpenStudy (anonymous):

Not entirely sure what you mean to be honest with you, haha. :P

OpenStudy (goformit100):

@mathslover @modphysnoob @Mathhater94 HELP in This question

OpenStudy (anonymous):

I got it figured out, thanks for the help. going to close the question.

OpenStudy (anonymous):

limits of integration were 7pi/6 and 11pi/6

mathslover (mathslover):

I think it would be \(\cfrac{7 \pi}{6} \) to \(\cfrac{3\pi}{2}\)

OpenStudy (anonymous):

3pi/2 is right on the y axis, would that still work?

mathslover (mathslover):

The curve reaches the highest point of the inner loop at \(\cfrac{3\pi}{2}\) . I had used symmetry here : the function can be evaluated from \(\cfrac{7\pi}{6}\) to \(\cfrac{3\pi}{2} \) and multiply by 2 to get the area of the entire loop.

mathslover (mathslover):

\(2\int _ {\cfrac{7\pi}{6}} ^{\cfrac{3\pi}{2}} \cfrac{1}{2} (1+2\sin \theta)^2 d\theta\) \(\implies \int _ {\cfrac{7\pi}{6}} ^{\cfrac{3\pi}{2}} 1 + 4\sin \theta +4(\cfrac{1-\cos 2\theta}{2})d\theta\) \(\implies [3\theta +4\cos \theta -sin 2\theta ]_{\cfrac{7\pi}{6}}^{\cfrac{3\pi}{2}}\) \(\textbf{which gives us} \space = \pi - \cfrac{3\sqrt{3}}{2}\)

mathslover (mathslover):

If I am right then ^ above will be the correct work.